# Quantum Data Management in the NISQ Era: Extended Version

### v2

@@ Abstract.
Quantum computing has emerged as a promising tool for transforming the landscape of computing technology. Recent efforts have applied quantum techniques to classical database challenges, such as query optimization, data integration, index selection, and transaction management. In this paper, we shift focus to a critical yet underexplored area: data management for quantum computing. We are currently in the Noisy Intermediate-Scale Quantum (NISQ) era, where qubits, while promising, are fragile and still limited in scale. After differentiating quantum data from classical data, we outline current and future data management paradigms in the NISQ era and beyond. We address the data management challenges arising from the emerging demands of near-term quantum computing. Our goal is to chart a clear course for future quantum-oriented data management research, establishing it as a cornerstone for the advancement of quantum computing in the NISQ era.
@@ 1. Introduction
While recent vision papers, tutorials, and surveys (Çalikyilmaz et al., 2023; Hai et al., 2024; Yuan et al., 2023, 2024; Winker et al., 2023b) have begun discussing the intersection of data management and quantum computing, they primarily focus on how quantum technologies can accelerate classical database operations. In contrast, our work delves into an equally important yet underexplored area: data management for quantum computing. At the moment, the fundamental concepts, research directions, and problem definitions in this area remain obscure within the database community. This paper initiates the exploration of these critical aspects and lays the foundation for further deeper integration of data management and quantum computing. We will not dive into the basics of quantum computing and information, as ample resources are available (Young et al., 2023; Quist et al., 2024; Xu et al., 2023; Biamonte, 2020). Instead, we focus on outlining the vision for data management paradigms in the NISQ era and identifying potential research challenges that are particularly relevant to the database community, which could have a significant impact on the advancement of today’s quantum technologies.
@@
  • Fundamentals: We explain quantum data, differentiate it from classical data (Sec. 2).
@@ 2. Data in the quantum era
1. Quantum data is probabilistic. Unlike a classical bit, which is 0 or 1, a quantum bit can be in superposition. Mathematically, a zero state and a one state may be represented by a unit vector in the standard basis. That is, the zero state |$\ket{0}$ is represented by the vector $\begin{bmatrix}1\\ 0\end{bmatrix}$ and the one state |$\ket{1}$ is represented by the vector $\begin{bmatrix}0\\ 1\end{bmatrix}$.  A single qubit state, denoted |$\ket{\psi}$, may be represented by a superposition, i.e., a linear combination of |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, for some pair of complex numbers $\alpha,\beta$ called amplitudes, which satisfy $|\alpha|^{2}+|\beta|^{2}=1$.  The probabilistic nature arises when a measurement is performed. When a measurement is performed on the state |$\ket{\psi}$, an outcome 0 is obtained with probability $|\alpha|^{2}$ and 1 obtained with probability $|\beta|^{2}$, and the state permanently changes to the obtained outcome. In classical computing, individual bits can be concatenated to form a bit string, e.g., the three-bit string “010.” Similarly, we may concatenate qubits into a multi-qubit state. In this case, an $n$-qubit state can be represented as a superposition of |$\ket{x}$ for $x\in\{0,1\}^{n}$, or equivalently a vector of $2^{n}$ components.
@@
|$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1\\ 0\\ 0\\ 1\end{smallmatrix}\right],$
where knowledge of one qubit determines the other. See Appendix A of our online report (Hai et al., 2025) for more background.
@@ 3. Our Vision: Data Management for Quantum Computing
This paradigm presents great potential for new database challenges. It focuses on using classical data to represent and simulate quantum data. This paradigm is more accessible as it relies on classical computers, not quantum ones.
@@
The potential for database research in this paradigm is immense. First, imagine innovative database systems specifically designed to manage classical data that supports simulation. Second, efficient caching strategies for expensive operations, such as repeated simulations of quantum circuits with varying error parameters, present another important research direction. The third research direction in this paradigm involves the representation of quantum states and operations. This is useful for quantum design questions such as optimizing the number of quantum gates or gate depth (Kissinger and van de Wetering, 2020; Zulehner et al., 2018) and compiling textbook quantum circuits to only include the low-level operations that real quantum devices support (Smith et al., 2020). As the representation of a quantum state (a vector) or quantum gate (a matrix) generally grows exponentially with the number of qubits, developing efficient methods to store and process these data structures is key to making simulations of quantum computation feasible, either in vector form or through more compressed structures like tensor networks (Biamonte, 2020; Orús, 2019) or other advanced representations (Bravyi and Gosset, 2016; van de Wetering, 2020a; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
@@
This paradigm focuses on the situation involved in most quantum-technology applications: classical computers handle the preprocessing and postprocessing of data for quantum devices, such as quantum chips, quantum sensors, and quantum computers. In this paradigm, the focus is on data which is primarily stored and processed as classical data. For instance, a classical computer may send instructions to a quantum chip. These instructions are classical data and, for example, describe the quantum circuit that should be run. The quantum device runs its computation and, after some time, performs measurements onto one or more of its qubits. The measurement outcomes, which are bits and hence classical data, are then returned to the classical computer. Thus, the quantum device receives classical input (the instructions) and returns the classical output (the measurement outcomes).
@@
Finally, we envision a future research paradigm beyond NISQ, when we have large-scale, fault-tolerant quantum computers. We will deal with pure quantum data, i.e., qubits. Here, the quantum hardware supports storing qubits for long enough to perform other tasks in the meantime. This enables for example a cloud quantum computer (Broadbent et al., 2009) which rotates its resources among multiple clients, various types of quantum machine learning (Cerezo et al., 2022), and Quantum Random Access Memory (Giovannetti et al., 2008), where data can be queried in superposition. A recent vision (Huang et al., 2022) shows a possible future: quantum data is collected from quantum sensing systems, e.g., for discovering a black hole; then stored and processed via the quantum memory of a quantum computer.
The data management research in this paradigm will be centered around handling quantum data, possibly developing quantum-native algorithms. Moreover, storing quantum data will remain challenging, as quantum data storage will still be costly in various ways (the number of qubits is not unbounded, robust storage of quantum data requires large-scale quantum error correction, which is computationally intensive, etc.). Another research topic is to efficiently allocate the available qubits. This includes, for example, various scheduling and allocation tasks (Chandra et al., 2022; Fittipaldi et al., 2023; Meter et al., 2009; Skrzypczyk and Wehner, 2021; Gauthier et al., 2023; Higgott and Breuckmann, 2021) and efficient use of different quantum hardware types each of which has speed-decoherence trade-offs (De Leon et al., 2021). Given the amount of quantum data in this paradigm, most research here will call for novel quantum data processing algorithms and systems beyond NISQ era.
@@ 4.1. The challenge of simulating quantum computation
The major practical bottleneck of (strong) simulation is scalability (Boixo et al., 2018; Li et al., 2020; Xu et al., 2023; Young et al., 2023). When simulating a quantum state of $n$ qubits, the state is typically represented as a vector with a size of $2^{n}$. The size of the state vector grows exponentially as $n$ increases. For instance, an experiment demonstrating quantum supremacy required 2.25 petabytes for 48 qubits, reaching the memory limits of today’s supercomputers (Boixo et al., 2018). To overcome the memory restriction, existing simulation tools have explored approximation (Vidal, 2003; Medvidović and Carleo, 2021; Jónsson et al., 2018), data compression (Wu et al., 2019), parallelization (Huang et al., 2021), and distributed computing (Smelyanskiy et al., 2016). In Fig. 2 we illustrate concepts mentioned in this section.
@@ 4.2. Databases to the rescue
At its core, a CQSS must, at a minimum, be capable of evaluating quantum circuits. This primarily involves performing linear algebraic operations, often described in terms of tensor networks (Biamonte, 2019). [^3] For simplicity of exposition, we do not consider non-tensor-based simulation methods which represent quantum states by their symmetries, rather than by complex-valued vectors (Gottesman, 1998). A tensor is a multidimensional array, and the dimensions along which a tensor extends are its indices. A tensor network consists of vertices representing tensors, and edges the indices. Free indices are depicted as legs, i.e., edges that only connect to one vertex in the graph. Connecting two vertices by joining a leg corresponds to the contraction with the corresponding indices. Fig. 2b shows an example of the tensor network representation of GHZ state. Tensor contraction is the summation of shared indices between tensors and is a generalization of matrix multiplication (Biamonte, 2020). See Appendix A of (Hai et al., 2025) for more details. The first question we ask is as follows.
@@
Figure 3. Memory usage comparison of three RDBMS-based solutions (red) against a general approach, i.e., order-n tensor baseline (blue) and the SotA MPS baseline (green). The x-axis shows increasing numbers of qubits, and the y-axis shows memory consumption (KB) in $\log_{10}$-scale. For sparse circuit (GHZ), RDBMS solutions demonstrate significantly lower memory usage than the order-n tensor baseline, with SQLite showing a slight advantage over the MPS baseline. For the dense circuit QFT, the memory usage of RDBMS solutions grows comparably to or exceeds the order-n tensor baseline, indicating the need for improvements in handling dense tensor computations.
@@
In other words, should we map the simulation workload to SQL queries and store the results in a DBMS? The advantage is that we can benefit from the efficiency and portability of modern database systems. And indeed, initial efforts from the database community (Blacher et al., 2023; Trummer, 2024) have begun to address the out-of-core simulation challenge by mapping quantum states and gates to SQL queries. For instance, in (Blacher et al., 2023), qubit states and gates are represented as tensor networks, with sparse tensors in COO format [^4] https://docs.scipy.org/doc/scipy/reference/generated/scipy.sparse.coo_matrix.html and Einstein summation operations mapped to SQL queries. Additionally, the recent abstract (Trummer, 2024) introduces the idea of supporting dense vectors using SQL UNION ALL and CASE statements. Despite these efforts that support basic quantum states and gates up to 18 qubits and a circuit depth of 18 in separate experiments, the core challenges of representing general quantum states and operations in DBMSs and achieving scalability remain. Moreover, our preliminary results in Fig. 3 demonstrate that RDBMS-based simulations are efficient for quantum circuits involving sparse tensor computations. All experimental details, together with a space and time complexity analysis, can be found in Appendix B of (Hai et al., 2025). RDBMS-based solutions are competitive with, and may even outperform, state-of-the-art representations like matrix product states (MPS) (Pérez-Garcia et al., 2007) in specific scenarios. Interestingly, this advantage does not extend to dense circuits such as quantum fourier transform (QFT) circuits, highlighting the need for further optimizations to enhance RDBMS systems for simulation workloads. In Appendix C of (Hai et al., 2025) we additionally discuss the use of NoSQL databases. On the other hand, the connection to tensor computations suggests the following question.
@@
[roundcorner=3pt,innerleftmargin = 5pt] Q2. Should we instead leverage tensor-based database technologies for simulation? Indeed, the database community has a substantial body of work on tensor computation, especially with the recent synergies between databases and machine learning (ML) such as learning ML models over relational data (Khamis et al., 2020; Boehm et al., 2023; Makrynioti and Vassalos, 2019; Zhou et al., 2020a; Paganelli et al., 2023). These studies range from high-level representations to the runtime optimization of ML workflows. At the representation level, numerous studies represent data as tensors (Koutsoukos et al., 2021) and aim to integrate relational and tensor operations (Schleich et al., 2016; Khamis et al., 2018; Huang et al., 2023). Meanwhile, system-building endeavors from the database community (Gao et al., 2017; Luo et al., 2017; Nikolic and Olteanu, 2018; Schleich et al., 2019; Boehm et al., 2020; Nakandala et al., 2020; He et al., 2022; Park et al., 2022; Boehm et al., 2020; Schleich et al., 2023) focus on performance efficiency and scalability, employing optimization techniques during compilation (e.g., operator fusion (Boehm et al., 2018, 2023)) and runtime (e.g., parallelization (Boehm et al., 2014)). Similar endeavours exist in the areas of compilers and High Performance Computing (Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Chen et al., 2018; Jia et al., 2019).
Our envisioned classical-quantum simulation system holds the potential to achieve optimal performance. By integrating relevant optimization techniques from the aforementioned existing systems, we can facilitate the automatic optimization of the underlying tensor network. The actual runtime performance will depend on several factors: the operations within the tensor network, the sparsity of the tensors involved, the memory layouts (i.e., how tensors are stored in memory, such as row-major or column-major formats), and the available hardware (Boehm et al., 2023; Schleich et al., 2023).
To address challenges unique to quantum computing, we highlight two key gaps that need to be overcome:

### v3

@@ Abstract.
Quantum computing has emerged as a transformative force in the evolution of computing technology. Recent efforts have applied quantum techniques to classical database challenges, such as query optimization, data integration, index selection, and transaction management. In this paper, we shift focus to a critical yet underexplored area: data management for quantum computing. We are currently in the noisy intermediate-scale quantum (NISQ) era, where qubits, while promising, are fragile and still limited in scale. After differentiating quantum data from classical data, we outline current and future data management paradigms in the NISQ era and beyond. We address the data management challenges arising from the emerging demands of near-term quantum computing. Our goal is to chart a clear course for future quantum-oriented data management research, establishing it as a cornerstone for the advancement of quantum computing in the NISQ era.
@@ 1. Introduction
While recent vision papers, tutorials, and surveys (Çalikyilmaz et al., 2023; Hai et al., 2024; Yuan et al., 2023, 2024; Winker et al., 2023b) have begun discussing the intersection of data management and quantum computing, they primarily focus on how quantum technologies can accelerate classical database operations. In contrast, our work delves into an equally important yet underexplored area: data management for quantum computing. At the moment, the fundamental concepts, research directions, and problem definitions in this area remain obscure within the database community. This paper initiates the exploration of these critical aspects and lays the foundation for further deeper integration of data management and quantum computing. We will not dive into the basics of quantum computing and information, as ample resources are available (Young et al., 2023; Quist et al., 2024; Xu et al., 2023; Biamonte, 2020). Instead, we focus on outlining the vision for data management paradigms in the NISQ era and on identifying potential research challenges that are particularly relevant to the database community, i.e., which could have a significant impact on the advancement of today’s quantum technologies.
@@
  • Fundamentals: We first explain quantum data and differentiate it from classical data (Sec. 2).
@@ 2. Data in the quantum era
1. Quantum data is probabilistic. Unlike a classical bit, which is 0 or 1, a quantum bit can be in superposition. Mathematically, a zero state and a one state may be represented by a unit vector in the standard basis. That is, the zero state |$\ket{0}$ is represented by the vector $\begin{bmatrix}1\\ 0\end{bmatrix}$ and the one state |$\ket{1}$ is represented by the vector $\begin{bmatrix}0\\ 1\end{bmatrix}$.  A single qubit state, denoted |$\ket{\psi}$, may be represented by a superposition, i.e., a linear combination of |$\ket{0}$ and |$\ket{1}$: |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, for some pair of complex numbers $\alpha,\beta$ called amplitudes, which satisfy $|\alpha|^{2}+|\beta|^{2}=1$.  The probabilistic nature arises when a measurement is performed. When a measurement is performed on the state |$\ket{\psi}$, the outcome 0 is obtained with probability $|\alpha|^{2}$ and the outcome 1 is obtained with probability $|\beta|^{2}$, and the state permanently changes to the obtained outcome. In classical computing, individual bits can be concatenated to form a bit string, e.g., the three-bit string “010.” Similarly, we may concatenate qubits into a multi-qubit state. In this case, an $n$-qubit state can be represented as a superposition of |$\ket{x}$ for $x\in\{0,1\}^{n}$, or equivalently a vector of $2^{n}$ components.
@@
|$\ket{\Psi}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})=\frac{1}{\sqrt{2}}\left[\begin{smallmatrix}1\\ 0\\ 0\\ 1\end{smallmatrix}\right],\vspace{-0.1cm}$
where knowledge of one qubit determines the other. Please see Appendix A of our online report (Hai et al., 2025) for more background.
@@ 3. Our Vision: Data Management for Quantum Computing
This paradigm presents great potential for new database challenges. It focuses on using classical data to represent and simulate quantum data. For example, Bell’s state from Sec. 2 is represented by a classical vector. This paradigm is more accessible as it relies on classical computers, not quantum ones.
@@
The potential for database research in this paradigm is immense. First, imagine innovative database systems specifically designed to manage classical data that supports simulation. Second, efficient caching strategies for expensive operations, such as repeated simulations of quantum circuits with varying error parameters, present another important research direction. The third research direction in this paradigm involves the representation of quantum states and operations. This is useful for quantum design questions such as optimizing the number of quantum gates or gate depth (Kissinger and van de Wetering, 2020; Zulehner et al., 2018) and compiling textbook quantum circuits to only include the low-level operations that real quantum devices support (Smith et al., 2020). As the representation of a quantum state (a vector) or quantum gate (a matrix) generally grows exponentially with the number of qubits, developing efficient methods to store and process these data structures is key to making simulations of quantum computation feasible, either in vector form or through more compressed structures like tensor networks (Biamonte, 2020; Orús, 2019) or other advanced representations (Bravyi and Gosset, 2016; van de Wetering, 2020; Zulehner and Wille, 2018; Jozsa and Miyake, 2008).
@@
This paradigm focuses on the situation involved in most quantum-technology applications: classical computers handle the preprocessing and postprocessing of data for quantum devices, such as quantum chips, quantum sensors, and quantum computers. In this paradigm, the focus is on data which is primarily stored and processed as classical data. For instance, a classical computer sends instructions (classical data) to a quantum chip, which performs computations and returns measurement outcomes—also classical data—back to the classical computer. Thus, the quantum device takes classical input and produces classical output.
@@
Finally, we envision a future research paradigm beyond NISQ, when we have large-scale, fault-tolerant quantum computers. We will deal with pure quantum data, i.e., qubits. Here, the quantum hardware supports storing the qubits for long enough to perform other tasks in the meantime. This enables for example a cloud quantum computer (Broadbent et al., 2009) that rotates its resources among multiple clients, various types of quantum machine learning (Cerezo et al., 2022), and quantum random access memory (Giovannetti et al., 2008), where data can be queried in superposition. A recent vision (Huang et al., 2022) shows a possible future: quantum data is collected from quantum sensing systems, e.g., for discovering a black hole; then stored and processed via the quantum memory of a quantum computer.
The data management research in this paradigm will be centered around handling quantum data, possibly developing quantum-native algorithms. Moreover, storing quantum data will remain challenging, as quantum data storage will still be costly in various ways (the number of qubits is not unbounded, robust storage of quantum data requires large-scale quantum error correction, which is computationally intensive, etc.). Another research topic is to efficiently allocate the available qubits. This includes, for example, various scheduling and allocation tasks (Chandra et al., 2022; Fittipaldi et al., 2023; Meter et al., 2009; Skrzypczyk and Wehner, 2021; Gauthier et al., 2023; Higgott and Breuckmann, 2021) and efficient use of different quantum hardware types each of which has speed-decoherence trade-offs (De Leon et al., 2021). Given the amount of quantum data in this paradigm, most research here calls for novel quantum data processing algorithms and systems beyond NISQ era.
@@ 4.1. The Challenge of Simulating Quantum Computation
The major practical bottleneck of (strong) simulation is scalability (Boixo et al., 2018; Li et al., 2020; Xu et al., 2023; Young et al., 2023). When simulating a quantum state of $n$ qubits, the state is typically represented as a vector with a size of $2^{n}$. The size of the state vector grows exponentially as $n$ increases. For instance, an experiment demonstrating quantum supremacy required 2.25 petabytes for 48 qubits, reaching the memory limits of today’s supercomputers (Boixo et al., 2018). To overcome the memory restriction, existing simulation tools have explored approximation (Vidal, 2003; Medvidović and Carleo, 2021; Jónsson et al., 2018), data compression (Wu et al., 2019), parallelization (Huang et al., 2021), and distributed computing (Smelyanskiy et al., [n.d.]). In Fig. 2 we illustrate concepts mentioned in this section.
@@ 4.2. Databases to the Rescue
At its core, a CQSS must, at a minimum, be capable of evaluating quantum circuits. This primarily involves performing linear algebraic operations, often described in terms of tensor networks (Biamonte, 2020). [^3] For simplicity of exposition, we do not consider non-tensor-based simulation methods which represent quantum states by their symmetries, rather than by complex-valued vectors (Gottesman, 1998). A tensor is a multidimensional array, and the dimensions along which a tensor extends are its indices. A tensor network consists of vertices representing tensors, and edges the indices. Free indices are depicted as legs, i.e., edges that only connect to one vertex in the graph. Connecting two vertices by joining a leg corresponds to the contraction with the corresponding indices. Fig. 2b shows the tensor network representation of the GHZ state. Tensor contraction is the summation of shared indices between tensors and is a generalization of matrix multiplication (Biamonte, 2020). See Appendix A of (Hai et al., 2025) for more details. The first question we ask is as follows.
@@
Figure 3. Memory usage comparison of three RDBMS-based solutions (red) against a general approach, i.e., order-n tensor baseline (blue) and the SotA MPS baseline (green), both implemented in Python. The x-axis shows increasing numbers of qubits, and the y-axis shows memory consumption (KB) in $\log_{10}$-scale. For sparse circuit (GHZ), RDBMS solutions demonstrate significantly lower memory usage than the order-n tensor baseline, with SQLite showing a slight advantage over the MPS baseline. For the dense circuit QFT, the memory usage of RDBMS solutions grows comparably to or exceeds the order-n tensor baseline, indicating the need for improvements in handling dense tensor computations.
[…]
@@
In other words, should we map the simulation workload to SQL queries and store the results in a DBMS? The advantage is that we can benefit from the efficiency and portability of modern database systems. And indeed, initial efforts from the database community (Blacher et al., 2023; Trummer, 2024) have begun to address the out-of-core simulation challenge by mapping quantum states and gates to SQL queries. For instance, in (Blacher et al., 2023), qubit states and gates are represented as tensor networks, with sparse tensors in COO format and Einstein summation operations mapped to SQL queries. Additionally, the recent abstract (Trummer, 2024) introduces the idea of supporting dense vectors using SQL UNION ALL and CASE statements. Despite these efforts that support basic quantum states and gates up to 18 qubits and a circuit depth of 18 in separate experiments, the core challenges of representing general quantum states and operations in DBMSs and achieving scalability remain. Moreover, our preliminary results in Fig. 3 demonstrate that RDBMS-based simulations are efficient for quantum circuits involving sparse tensor computations. All experimental details, together with a space and time complexity analysis, can be found in Appendix B of (Hai et al., 2025). An early prototype is described in (Littau and Hai, 2025). RDBMS-based solutions are competitive with, and may even outperform, state-of-the-art representations like matrix product states (MPS) (Pérez-Garcia et al., 2007) in specific scenarios. Interestingly, this advantage does not extend to dense circuits such as quantum fourier transform (QFT) circuits, highlighting the need for further optimizations to enhance RDBMS systems for simulation workloads. In Appendix C of (Hai et al., 2025) we also discuss the use of NoSQL databases.
Q1 summary. Pushing simulation workloads to DBMSs leverages their efficiency, showing promise for sparse circuits, yet challenges in optimization and scalability remain.
On the other hand, the connection to tensor computations suggests the following question.
@@
[roundcorner=3pt,innerleftmargin = 5pt] Q2. Should we instead leverage tensor-based database technologies for simulation? Indeed, the database community has a substantial body of work on tensor computation, especially with the recent synergies between databases and machine learning (ML) such as learning ML models over relational data (Khamis et al., 2020; Boehm et al., 2023; Makrynioti and Vassalos, 2019; Zhou et al., 2020a; Paganelli et al., 2023). These studies range from high-level representations to the runtime optimization of ML workflows. At the representation level, numerous studies represent data as tensors (Koutsoukos et al., 2021) and aim to integrate relational and tensor operations (Schleich et al., 2016; Khamis et al., 2018; Huang et al., 2023). Meanwhile, system-building efforts from the database community (Gao et al., 2017; Luo et al., 2017; Nikolic and Olteanu, 2018; Schleich et al., 2019; Boehm et al., 2020; Nakandala et al., 2020; He et al., 2022; Park et al., 2022; Boehm et al., 2020; Schleich et al., 2023) focus on performance efficiency and scalability, employing optimization techniques during compilation (e.g., operator fusion (Boehm et al., 2018, 2023)) and runtime (e.g., parallelization (Boehm et al., 2014)). Similar endeavours exist in the areas of compilers and high performance computing (HPC) (Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Chen et al., 2018; Jia et al., 2019).
Our envisioned classical-quantum simulation system holds the potential to achieve optimal performance. By integrating relevant optimization techniques from the aforementioned existing systems, we can facilitate the automatic optimization of the underlying tensor network. The actual runtime performance will depend on several factors: the operations within the tensor network, the sparsity of the tensors involved, the memory layouts (i.e., how tensors are stored in memory, such as row-major or column-major formats), and the available hardware (Boehm et al., 2023; Schleich et al., 2023). To address challenges unique to quantum computing, we highlight two key gaps below.